# pFedMMA: Personalized Federated Fine-Tuning with Multi-Modal Adapter for Vision-Language Models

### v2

@@ Abstract
Vision-Language Models (VLMs) like CLIP have demonstrated remarkable generalization in zero- and few-shot settings, but adapting them efficiently to decentralized, heterogeneous data remains a challenge. While prompt tuning has emerged as a popular parameter-efficient approach in personalized federated learning, existing methods often sacrifice generalization in favor of personalization, struggling particularly on unseen classes or domains. In this work, we propose pFedMMA, the first personalized federated learning framework that leverages multi-modal adapters for vision-language tasks. Each adapter contains modality-specific up- and down-projection layers alongside a globally shared projection that aligns cross-modal features. Our optimization strategy allows clients to locally adapt to personalized data distributions while collaboratively training the shared projection to improve global generalization. This design is also communication-efficient, as only the shared component is exchanged during communication rounds. Through extensive experiments across eleven datasets, including domain- and label-shift scenarios, we show that pFedMMA achieves state-of-the-art trade-offs between personalization and generalization, outperforming recent federated prompt tuning methods.
@@ 1 Introduction
Vision-Language Models (VLMs) like CLIP Radford et al. (2021) have revolutionized multi-modal learning by jointly embedding visual and textual data through massive contrastive pre-training Jia et al. (2021); Li et al. (2022); Yao et al. . This paradigm empowers models to generalize effectively in zero-shot and few-shot settings Zhang et al. (2022); Zhu et al. (2023); Ghiasvand et al. (2025); Aghdam & Hu (2025). Among them, larger transformer-based variants Vaswani (2017) (e.g., CLIP ViT-L/14) consistently outperform smaller counterparts such as ViT-B/16, with margins exceeding 6% on benchmarks like ImageNet Deng et al. (2009). However, the computational demands of fine-tuning such large-scale models with billions of parameters pose significant challenges, particularly for domain-specific tasks Oskouie et al. (2025). To mitigate this, Parameter-Efficient Fine-Tuning (PEFT) techniques have emerged, especially in NLP. These methods, including adapters Chen et al. (2022); Karimi Mahabadi et al. (2021); Rebuffi et al. (2017) and prompt tuning Jia et al. (2022); Li & Liang (2021), introduce a lightweight set of trainable parameters or tokens, allowing the backbone model to remain frozen.
@@
While highly effective in centralized settings, these techniques fall short in scenarios involving decentralized and privacy-sensitive data, such as healthcare, legal, or industrial domains Manoel et al. (2023); Shoham & Rappoport (2023); Mahjourian & Nguyen (2025). Federated Learning (FL) offers a promising alternative by enabling collaborative training without raw data sharing. In FL, clients update their local models and transmit only intermediate model updates such as parameters or gradients, which are aggregated into a global model by a central server McMahan et al. (2017).
@@
In real-world scenarios, client data often exhibits variations in domain discrepancies (feature shift) Li et al. or imbalanced class distributions (label shift) Li et al. (2021a). Simply applying standard aggregation strategies, such as FedAvg McMahan et al. (2017), over prompts Guo et al. (2023b) or other fine-tuning methods, such as LoRA, often leads to suboptimal performance due to data heterogeneity Zhang et al. (2023); Borazjani et al. (2025). As a result, Personalized Federated Learning (PFL), particularly with prompt tuning, has gained increasing attention. pFedPrompt Guo et al. (2023a) introduces personalization by coupling a global text prompt with local visual attention modules to tailor predictions to each client’s data. FedOTP Li et al. (2024) uses Optimal Transport to align local and global representations under label shift. FedPGP Cui et al. applies prompt-wise contrastive learning to enhance inter-client generalization. Recently, pFedMoAP Luo et al. (2025) proposes a Mixture-of-Experts framework, where prompts from other clients serve as non-local experts, and each client learns an attention-based gating mechanism for selective adaptation. While these methods achieve impressive personalization performance, they often struggle to generalize to unseen classes or domains, limiting their applicability in out-of-distribution scenarios. For example, as shown in Fig. 1, FedOTP achieves poor harmonic mean accuracy, even though it has been shown to have strong personalization performance.
@@
Beyond prompt tuning, adapters offer another PEFT strategy by introducing small trainable modules into frozen pre-trained models Cai et al. (2020); Chen et al. (2022; b); Gao et al. (2024); Hu et al. (2021); Zhang et al. (2022). Unlike prompts, adapters operate independently of model architecture and can be easily inserted into various backbones, such as ResNets He et al. (2016), ViTs Dosovitskiy et al. , and Swin Transformers Liu et al. (2021). However, most adapter methods like AdaptFormer Chen et al. (2022) and LoRA Hu et al. (2021) are uni-modal and do not account for the cross-modal dependencies inherent in VLMs like CLIP Radford et al. (2021). Multi-modal adapters Yang et al. (2024) address this by integrating both visual and textual signals via a shared projection layer that promotes feature alignment across modalities while preserving modality-specific knowledge. Despite their demonstrated advantages over prompt-based approaches Yang et al. (2024); Guo & Gu (2025), their integration with PFL remains largely unexplored.
@@
Before delving into details, we summarize our contributions: (1) We observe that while most state-of-the-art prompt tuning methods achieve strong personalization performance, they often generalize poorly to unseen classes. To address this, we introduce a multi-modal adapter framework that explicitly aims to balance personalization and generalization in federated vision-language learning. (2) We propose pFedMMA, the first adapter-based approach for PFL of VLMs. Our architecture incorporates modality-specific up- and down-projection layers and a shared cross-modal projection. All components are updated locally, but only the shared projection is aggregated globally, enabling effective asymmetric optimization. (3) We conduct extensive experiments on widely used benchmarks to evaluate pFedMMA’s performance on base-to-novel generalization across both category- and domain-level tasks under heterogeneous data distributions. Results demonstrate the superiority of our approach in harmonizing generalization and personalization.
@@ 2.2 Vision-Language Classification with Few-Shot Adaptation
This fine-tuning step enables the model to better capture domain-specific semantics while maintaining the efficiency and generalization capabilities of the pretrained architecture. Adaptation can be achieved through various strategies. One approach is to directly optimize the input prompts $\{c_{k}\}_{k=1}^{K}$, following the principles of prompt tuning Chen et al. (a). Alternatively, lightweight task-specific modules such as adapter layers Gao et al. (2024) or low-rank parameterizations like LoRA Zanella & Ben Ayed (2024) can be fine-tuned, while keeping the backbone encoders frozen.
@@ 3.1 Multi-Modal Adapter Architecture
The motivation for this design stems from two empirical findings: (i) higher layers of both image and text encoders in CLIP contain more discriminative and dataset-specific features, while lower layers preserve general, transferable knowledge; and (ii) larger modality gaps between text and image encoders are observed in the lower layers, making cross-modal alignment particularly challenging in the early stages of the network. Yang et al. (2024).
@@
In contrast to methods that inject prompts or adapters across all layers Chen et al. (2022); Houlsby et al. (2019); Hu et al. (2021) or some lower layers Khattak et al. (2023a; b); Zhou et al. (2022b; a), this selective, top-layer insertion strategy reduces the number of trainable parameters while maintaining cross-modal adaptability.
@@ 3.2 Generalization and Personalization via pFedMMA
After local updates, only the shared projection parameters ${\bm{W}}_{js,i}^{t,E}$ are uploaded to the server. These are aggregated across all participating clients to obtain the updated global shared adapter: ${\bm{W}}_{js}^{t+1}=\frac{1}{N}\sum_{i=1}^{N}{\bm{W}}_{js,i}^{t,E},$ where $N$ is the number of participating clients. In contrast, the up- and down-projection parameters remain local and are not shared or averaged.
@@ 4.1 Experimental Setup
Datasets and Data Heterogeneity. To evaluate the effectiveness of pFedMMA, we conduct experiments across eleven public benchmark datasets that cover various types of data heterogeneity, including label shift and feature shift. Following prior work such as Guo et al. (2023b), we use seven visual classification datasets: SUN397 Xiao et al. (2010), OxfordPets Parkhi et al. (2012), Flowers102 Nilsback & Zisserman (2008), DTD Cimpoi et al. (2014), Caltech101 Fei-Fei et al. (2004), UCF101 Soomro et al. (2012), and Food101 Bossard et al. (2014). We refer to these as the CLIP datasets. To simulate severe label heterogeneity, we apply a pathological non-IID setting in which each client is assigned a distinct, non-overlapping set of classes. Clients are trained on their local classes and evaluated on their own classes, on base classes held by other clients, and on novel classes that are unseen during the training process.
@@
Baselines. We evaluate pFedMMA across all experimental settings, including generalization, personalization, and domain generalization, using a consistent set of five baselines. Zero-shot CLIP Radford et al. (2021) serves as a non-adaptive reference model that uses fixed hand-crafted prompt templates such as ”a photo of a [class]” without any task-specific learning. PromptFL Guo et al. (2023b) represents a standard federated prompt learning approach in which a shared prompt is collaboratively learned across clients using FedAvg. FedPGP Cui et al. introduces prompt-wise contrastive learning to encourage consistency between global and local prompts. FedOTP Li et al. (2024) applies unbalanced Optimal Transport to align global knowledge with client-specific prompt representations. Finally, pFedMoAP Luo et al. (2025) leverages a Mixture-of-Experts design that enables each client to access both local and non-local prompt experts through a lightweight attention-based gating mechanism. These baselines cover a diverse range of federated adaptation strategies, providing a strong benchmark for assessing the performance of pFedMMA across different types of heterogeneity.
@@
PromptFL Guo et al. (2023b) 88.93 88.95 75.36 83.09 $77.73$ $77.71$ $72.96$ $76.07$ 97.37 97.06 63.62 82.66 $80.23$ $80.21$ $45.29$ $63.81$
FedPGP Cui et al. 95.38 76.49 71.68 79.09 $94.29$ $54.88$ $57.76$ $65.02$ 99.67 72.44 58.65 73.37 $89.03$ $71.03$ $50.94$ $66.75$
@@
PromptFL Guo et al. (2023b) $96.35$ $96.28$ $97.26$ $96.63$ $97.77$ $98.19$ $92.58$ $96.11$ $90.48$ $90.50$ $91.37$ $90.78$ $82.57$ $82.73$ $64.47$ $75.55$
FedPGP Cui et al. $96.62$ $95.17$ $97.15$ $96.31$ $99.42$ $94.94$ $90.88$ $94.95$ $93.70$ $86.38$ $87.14$ $88.96$ $94.94$ $60.62$ $59.26$ $68.33$
@@
Implimentation Details. All methods, including pFedMMA and all baselines, are implemented on top of a frozen CLIP model. We use two backbone architectures, ViT-B16 and ViT-B32 Dosovitskiy et al. , and default to ViT-B16 unless otherwise specified. For the CLIP datasets, each is split into $10$ clients with non-overlapping classes, using $100$ percent participation, $2$ local epochs, and $50$ communication rounds. For the CIFAR-10 and CIFAR-100 datasets, we simulate a large-scale federated environment with $100$ clients, using a varying Dirichlet distribution and a $10$ percent client participation rate per communication round. Training runs for $50$ rounds with 1 local epoch per round. In the case of DomainNet and Office-Caltech10, each domain of these two datasets is partitioned to $1/2$ clients, resulting in $N=6/12$ for DomainNet and $N=4/8$ for Office-Caltech10. We use SGD with a learning rate of $0.001$, and batch sizes of $32$ for training and $100$ for testing. Further implementation details are provided in Appendix.
@@ 4.2 Performance Evaluation
Base-to-Novel Class Generalization. We evaluate the performance of pFedMMA in terms of its ability to generalize from locally trained classes to both base and novel classes. Following prior work, we report top-1 accuracy on each client’s local classes, on the base classes seen by other clients, and on novel classes that are entirely unseen during training. To capture overall effectiveness, we use the harmonic mean (HM) of these three metrics. As summarized in Table 1 for the 16-shot setting, pFedMMA consistently achieves strong performance across all evaluation categories and delivers the best overall HM averaged across seven datasets, outperforming all baselines.
Zero-shot CLIP and PromptFL suffer from poor local accuracy, tending to favor generalization at the expense of personalization. We also report $\Delta$, which denotes the relative improvement of pFedMMA compared with the strongest non-baseline methods (FedPGP, FedOTP, and pFedMoAP). While FedOTP sometimes achieves high local accuracy, its extremely low base and novel class scores indicate poor generalization. pFedMoAP performs well on local classes due to its MoE-based prompt sharing, but it lags behind pFedMMA in base and novel accuracy. By contrast, pFedMMA achieves the highest base and novel accuracy, surpassing FedPGP and demonstrating excellent generalization, while remaining competitive on local classes—only $0.74\%$ lower than pFedMoAP.
@@
Model Evaluation on Feature & Label Shifts. To evaluate the robustness of pFedMMA in realistic federated learning scenarios, we examine its performance under both label shift and feature shift using the DomainNet and Office-Caltech10 datasets. Following the standard protocol, each domain is split into two clients via a Dirichlet distribution with $\beta=0.5$, yielding 12 clients for DomainNet and 8 clients for Office-Caltech10. The results in Table 3 show that under these challenging heterogeneous conditions, traditional methods such as CLIP and PromptFL struggle to generalize effectively. In contrast, pFedMMA consistently achieves the highest average accuracy across both datasets, highlighting its strong robustness to cross-domain shifts. Additional experiments with one or two clients per domain and varying $\beta$ are provided in Tables 9 and 10 in the Appendix.
4.3 Ablation Study

### v3

# arXiv:2507.05394v3[cs.CV] 27 Feb 2026

@@ Abstract
Vision-Language Models (VLMs) like CLIP have demonstrated remarkable generalization in zero- and few-shot settings, but adapting them efficiently to decentralized, heterogeneous data remains a challenge. While prompt tuning has emerged as a popular parameter-efficient approach in personalized federated learning, existing methods often sacrifice generalization in favor of personalization, struggling particularly on unseen classes or domains. In this work, we propose pFedMMA, a personalized federated learning framework that leverages multi-modal adapters for vision-language tasks. Each adapter contains modality-specific up- and down-projection layers alongside a globally shared projection that aligns cross-modal features. Our optimization strategy allows clients to locally adapt to personalized data distributions while collaboratively training the shared projection to improve global generalization. This design is also communication-efficient, as only the shared component is exchanged during communication rounds. Through extensive experiments across eleven datasets, including domain- and label-shift scenarios, we show that pFedMMA achieves state-of-the-art trade-offs between personalization and generalization, outperforming recent federated prompt tuning methods. Code is available at https://github.com/sajjad-ucsb/pFedMMA.
@@ 1 Introduction
Vision-Language Models (VLMs) like CLIP Radford et al. (2021) have revolutionized multi-modal learning by jointly embedding visual and textual data through massive contrastive pre-training Jia et al. (2021); Li et al. (2022); Yao et al. . This paradigm empowers models to generalize effectively in zero-shot and few-shot settings Zhang et al. (2022); Zhu et al. (2023); Ghiasvand et al. (2025a); Aghdam and Hu (2025). Among them, larger transformer-based variants Vaswani (2017) (e.g., CLIP ViT-L/14) consistently outperform smaller counterparts such as ViT-B/16, with margins exceeding 6% on benchmarks like ImageNet Deng et al. (2009). However, the computational demands of fine-tuning such large-scale models with billions of parameters pose significant challenges, particularly for domain-specific tasks Oskouie et al. (2025). To mitigate this, Parameter-Efficient Fine-Tuning (PEFT) techniques have emerged, especially in NLP. These methods, including adapters Chen et al. (2022); Karimi Mahabadi et al. (2021); Rebuffi et al. (2017) and prompt tuning Jia et al. (2022); Li and Liang (2021), introduce a lightweight set of trainable parameters or tokens, allowing the backbone model to remain frozen.
@@
While highly effective in centralized settings, these techniques fall short in scenarios involving decentralized and privacy-sensitive data, such as healthcare, legal, or industrial domains Manoel et al. (2023); Shoham and Rappoport (2023); Mahjourian and Nguyen (2025). Federated Learning (FL) offers a promising alternative by enabling collaborative training without raw data sharing. In FL, clients update their local models and transmit only intermediate model updates such as parameters or gradients, which are aggregated into a global model by a central server McMahan et al. (2017).
@@
In real-world scenarios, client data often exhibits variations in domain discrepancies (feature shift) Li et al. or imbalanced class distributions (label shift) Li et al. (2021a). Simply applying standard aggregation strategies, such as FedAvg McMahan et al. (2017), over prompts Guo et al. (2023a) or other fine-tuning methods, such as LoRA, often leads to suboptimal performance due to data heterogeneity Zhang et al. (2023); Borazjani et al. (2025). As a result, Personalized Federated Learning (PFL), particularly with prompt tuning, has gained increasing attention. pFedPrompt Guo et al. (2023b) introduces personalization by coupling a global text prompt with local visual attention modules to tailor predictions to each client’s data. FedOTP Li et al. (2024) uses Optimal Transport to align local and global representations under label shift. FedPGP Cui et al. (2024) applies prompt-wise contrastive learning to enhance inter-client generalization. Recently, pFedMoAP Luo et al. (2025) proposes a Mixture-of-Experts framework, where prompts from other clients serve as non-local experts, and each client learns an attention-based gating mechanism for selective adaptation. While these methods achieve impressive personalization performance, they often struggle to generalize to unseen classes or domains, limiting their applicability in out-of-distribution scenarios. For example, as shown in Fig. 1, FedOTP achieves poor harmonic mean accuracy, even though it has been shown to have strong personalization performance.
@@
Beyond prompt tuning, adapters offer another PEFT strategy by introducing small trainable modules into frozen pre-trained models Cai et al. (2020); Chen et al. (2022; ); Gao et al. (2024); Hu et al. (2021); Zhang et al. (2022). Unlike prompts, adapters operate independently of model architecture and can be easily inserted into various backbones, such as ResNets He et al. (2016), ViTs Dosovitskiy et al. , and Swin Transformers Liu et al. (2021). However, most adapter methods like AdaptFormer Chen et al. (2022) and LoRA Hu et al. (2021) are uni-modal and do not account for the cross-modal dependencies inherent in VLMs like CLIP Radford et al. (2021). Multi-modal adapters Yang et al. (2024) address this by integrating both visual and textual signals via a shared projection layer that promotes feature alignment across modalities while preserving modality-specific knowledge. Despite their demonstrated advantages over prompt-based approaches Yang et al. (2024); Guo and Gu (2025), their integration with PFL remains largely unexplored.
@@
Before delving into details, we summarize our contributions: (1) We observe that while most state-of-the-art prompt tuning methods achieve strong personalization performance, they often generalize poorly to unseen classes. To address this, we introduce a multi-modal adapter framework that explicitly aims to balance personalization and generalization in federated vision-language learning. (2) We propose pFedMMA, an adapter-based approach for PFL of VLMs. Our architecture incorporates modality-specific up- and down-projection layers and a shared cross-modal projection. All components are updated locally, but only the shared projection is aggregated globally, enabling effective asymmetric optimization. (3) We conduct extensive experiments on widely used benchmarks to evaluate pFedMMA’s performance on base-to-novel generalization across both category- and domain-level tasks under heterogeneous data distributions. Results demonstrate the superiority of our approach in harmonizing generalization and personalization.
@@ 2.2 Vision-Language Classification with Few-Shot Adaptation
This fine-tuning step enables the model to better capture domain-specific semantics while maintaining the efficiency and generalization capabilities of the pretrained architecture. Adaptation can be achieved through various strategies. One approach is to directly optimize the input prompts $\{c_{k}\}_{k=1}^{K}$, following the principles of prompt tuning Chen et al. . Alternatively, lightweight task-specific modules such as adapter layers Gao et al. (2024) or low-rank parameterizations like LoRA Zanella and Ben Ayed (2024) can be fine-tuned, while keeping the backbone encoders frozen.
@@ 3.1 Multi-Modal Adapter Architecture
The motivation for this design stems from two empirical findings: (i) higher layers of both image and text encoders in CLIP contain more discriminative and dataset-specific features, while lower layers preserve general, transferable knowledge; and (ii) larger modality gaps between text and image encoders are observed in the lower layers, making cross-modal alignment particularly challenging in the early stages of the network Yang et al. (2024).
@@
In contrast to methods that inject prompts or adapters across all layers Chen et al. (2022); Houlsby et al. (2019); Hu et al. (2021); Ghiasvand et al. (2025b) or some lower layers Khattak et al. (2023a; b); Zhou et al. (2022b; a), this selective, top-layer insertion strategy reduces the number of trainable parameters while maintaining cross-modal adaptability.
@@ 3.2 Generalization and Personalization via pFedMMA
After local updates, only the shared projection parameters ${\bm{W}}_{js,i}^{t,E}$ are uploaded to the server. These are aggregated across all participating clients to obtain the updated global shared adapter: ${\bm{W}}_{js}^{t+1}=\sum_{i=1}^{N}p_{i}{\bm{W}}_{js,i}^{t,E},$ where $p_{i}=\frac{n_{i}}{n}$ scales the contribution of each client by its dataset size $n_{i}$, with $n=\sum_{i}n_{i}$ and $N$ is the number of participating clients. In contrast, the up- and down-projection parameters remain local and are not shared or averaged.
@@ 4.1 Experimental Setup
Datasets and Data Heterogeneity. To evaluate the effectiveness of pFedMMA, we conduct experiments across eleven public benchmark datasets that cover various types of data heterogeneity, including label shift and feature shift. Following prior work such as Guo et al. (2023a), we use seven visual classification datasets: SUN397 Xiao et al. (2010), OxfordPets Parkhi et al. (2012), Flowers102 Nilsback and Zisserman (2008), DTD Cimpoi et al. (2014), Caltech101 Fei-Fei et al. (2004), UCF101 Soomro et al. (2012), and Food101 Bossard et al. (2014). We refer to these as the CLIP datasets. To simulate severe label heterogeneity, we apply a pathological non-IID setting in which each client is assigned a distinct, non-overlapping set of classes. Clients are trained on their local classes and evaluated on their own classes, on base classes held by other clients, and on novel classes that are unseen during the training process.
@@
Baselines. We evaluate pFedMMA across all experimental settings, including generalization, personalization, and domain generalization, using a consistent set of five baselines. Zero-shot CLIP Radford et al. (2021) serves as a non-adaptive reference model that uses fixed hand-crafted prompt templates such as “a photo of a [class]” without any task-specific learning. PromptFL Guo et al. (2023a) represents a standard federated prompt learning approach in which a shared prompt is collaboratively learned across clients using FedAvg. FedPGP Cui et al. (2024) introduces prompt-wise contrastive learning to encourage consistency between global and local prompts. FedOTP Li et al. (2024) applies unbalanced Optimal Transport to align global knowledge with client-specific prompt representations. Finally, pFedMoAP Luo et al. (2025) leverages a Mixture-of-Experts design that enables each client to access both local and non-local prompt experts through an attention-based gating mechanism. In addition to these prompt-based methods, we also consider adapter and LoRA-style PEFT baselines by implementing federated CLIP-Adapter Gao et al. (2024) and federated CLIP-LoRA Zanella and Ben Ayed (2024), where only the adapter or low-rank layers are updated and aggregated across clients. These baselines cover a diverse range of federated adaptation strategies, providing a strong benchmark for assessing the performance of pFedMMA across different types of heterogeneity.
@@
PromptFL Guo et al. (2023a) 88.93 $88.95$ $75.36$ $83.09$ $77.73$ $77.71$ $72.96$ $76.07$ $97.37$ $97.06$ $63.62$ $82.66$ $80.23$ $80.21$ $45.29$ $63.81$
FedCLIP-Adapter Gao et al. (2024) $78.97$ $79.22$ $82.09$ $80.04$ $72.82$ $72.78$ $77.22$ $74.22$ $71.46$ $73.22$ $77.38$ $73.94$ $56.99$ $55.90$ $59.78$ $57.51$
FedCLIP-LoRA Zanella and Ben Ayed (2024) $81.69$ $90.16$ $79.76$ $83.35$ $74.67$ $80.46$ $72.28$ $75.65$ $76.77$ $97.72$ $68.16$ $79.09$ $68.16$ $81.94$ $59.42$ $68.64$
FedPGP Cui et al. (2024) $95.38$ $76.49$ $71.68$ $79.09$ $94.29$ $54.88$ $57.76$ $65.02$ $99.67$ $72.44$ $58.65$ $73.37$ $89.03$ $71.03$ $50.94$ $66.75$
@@
PromptFL Guo et al. (2023a) $96.35$ $96.28$ $97.26$ $96.63$ $97.77$ $98.19$ $92.58$ $96.11$ $90.48$ $90.50$ $91.37$ $90.78$ $82.57$ $82.73$ $64.47$ $75.55$
FedCLIP-Adapter Gao et al. (2024) $93.01$ $92.93$ $97.09$ $94.30$ $96.28$ $97.35$ $94.10$ $95.89$ $90.07$ $90.10$ $91.19$ $90.45$ $72.13$ $72.23$ $77.88$ $73.98$
FedCLIP-LoRA Zanella and Ben Ayed (2024) $90.71$ $95.64$ $97.71$ $94.59$ $96.76$ $97.93$ $94.43$ $96.35$ $89.18$ $90.32$ $91.49$ $90.32$ $75.58$ $87.13$ $74.80$ $78.79$
FedPGP Cui et al. (2024) $96.62$ $95.17$ $97.15$ $96.31$ $99.42$ $94.94$ $90.88$ $94.95$ $93.70$ $86.38$ $87.14$ $88.96$ $94.94$ $60.62$ $59.26$ $68.33$
@@
Implimentation Details. All methods, including pFedMMA and all baselines, are implemented on top of a frozen CLIP model. We use two backbone architectures, ViT-B16 and ViT-B32 Dosovitskiy et al. , and default to ViT-B16 unless otherwise specified. For the CLIP datasets, each is split into $10$ clients with non-overlapping classes, using $100$ percent participation, $2$ local epochs, and $50$ communication rounds. For the CIFAR-10 and CIFAR-100 datasets, we simulate a large-scale federated environment with $100$ clients, using a varying Dirichlet distribution and a $10$ percent client participation rate per communication round. Training runs for $50$ rounds with 1 local epoch per round. In the case of DomainNet and Office-Caltech10, each domain of these two datasets is partitioned to $1/2$ clients, resulting in $N=6/12$ for DomainNet and $N=4/8$ for Office-Caltech10. We use SGD with a learning rate of $0.001$, and batch sizes of $32$ for training and $100$ for testing. Further implementation details are provided in the Appendix, where we also report additional results using the AdamW optimizer (Table 19), which exhibit similar trends to those obtained with SGD.
@@ 4.2 Performance Evaluation
Base-to-Novel Class Generalization.
We evaluate the performance of pFedMMA in terms of its ability to generalize from locally trained classes to both base and novel classes. Following prior work, we report top-1 accuracy on each client’s local classes, on the base classes seen by other clients, and on novel classes that are entirely unseen during training. To capture overall effectiveness, we use the harmonic mean (HM) of these three metrics, $\mathrm{HM}=3/(\mathrm{Acc}_{\text{local}}^{-1}+\mathrm{Acc}_{\text{base}}^{-1}+\mathrm{Acc}_{\text{novel}}^{-1}),$ which penalizes methods that over-optimize one component at the expense of the others and thus better reflects the balance between personalization (local) and generalization (base and novel) than a simple arithmetic mean; this type of harmonic-mean score is standard in generalized zero-shot learning and base-to-novel CLIP adaptation, and has also been adopted in recent PFL work to jointly summarize local, base, and novel accuracies Verma et al. (2020); Du and others (2025); Cui et al. (2024). As summarized in Table 1 for the 16-shot setting, pFedMMA consistently achieves strong performance across all evaluation categories and delivers the best overall HM averaged across seven datasets, outperforming all baselines.
Zero-shot CLIP, PromptFL, federated CLIP-Adapter, and CLIP-LoRA suffer from poor local accuracy, tending to favor generalization at the expense of personalization. We also report $\Delta$, which denotes the relative improvement of pFedMMA compared with the strongest non-baseline methods (FedPGP, FedOTP, and pFedMoAP). While FedOTP sometimes achieves high local accuracy, its extremely low base and novel class scores indicate poor generalization. pFedMoAP performs well on local classes due to its MoE-based prompt sharing, but it lags behind pFedMMA in base and novel accuracy. By contrast, pFedMMA achieves the highest base and novel accuracy, surpassing FedPGP and demonstrating excellent generalization, while remaining competitive on local classes—only $0.74\%$ lower than pFedMoAP.
@@
Model Evaluation on Feature & Label Shifts. To evaluate the robustness of pFedMMA in realistic federated learning scenarios, we examine its performance under both label shift and feature shift using the DomainNet and Office-Caltech10 datasets. Following the standard protocol, each domain is split into two clients via a Dirichlet distribution with $\beta=0.5$, yielding 12 clients for DomainNet and 8 clients for Office-Caltech10. The results in Table 3 show that under these challenging heterogeneous conditions, traditional methods such as CLIP and PromptFL struggle to generalize effectively. In contrast, pFedMMA consistently achieves the highest average accuracy across both datasets, highlighting its strong robustness to cross-domain shifts. Additional experiments with one or two clients per domain and varying $\beta$ are provided in Tables 11 and 12 in the Appendix.
4.3 Ablation Study & Hyperparameter Sensitivity